OpenStudy (anonymous):

What is 7/8 % of 24?

OpenStudy (anonymous):

A. 0.21 B. 2.1 C. 21 D. 210

OpenStudy (anonymous):

100% = 24 1% = 8/8% =? 1/8%=?